OpenStudy (anonymous):

estimate the sum or difference 1&5/6+2&1/2

OpenStudy (amistre64):

1 +5/6 +2 +1/2 = n 1(6) +5(6)/6 +2(6) +1(6)/2 = (6)n 6 +5 +12 +3 = 6n 24 = 6n 4

OpenStudy (amistre64):

err... 26/6 :) 4 1/3